# New theoretical Fe II templates for bright quasars

### v1

@@ 5 Summary
One of the most prominent characteristic features of the quasar spectrum is the iron emission. In this work, we examine the properties of Fe II emission using CLOUDY version C23.00 and generate new theoretical Fe II templates that can be used to model Fe II emission in the quasar spectra covering the wavelength range 1000-10000 Å. We study the impact of microturbulence and metallicity on iron emission and compare our model predictions with observational data. The primary outcomes of our study are outlined below.
@@
The long-standing problem of a mismatch between the expected and observed values of Fe II UV to optical ratio under the single zone approximation could not be resolved by the updated atomic database available in CLOUDY C23.00.
@@
4.
The ionisation flux has a more significant impact on the ratio of red-to-blue parts of optical Fe II around H than does the gas density.
5.
High-density BLR ( (cm) = 12 - 13) and extremely high radiation fluxes are needed for the highly accreting (xA) sources.
@@
7.
The observed UV-to-optical spectra of a quasar cannot be well fitted by a single theoretical Fe II template generated assuming fixed column density and isotropic cloud emission.

### v2

@@ 5 Summary
One of the most prominent characteristic features of the quasar spectrum is the iron emission. In this work, we examine the properties of Fe II emission using CLOUDY version C23.00 and generate new theoretical Fe II templates that can be used to model Fe II emission in the quasar spectra covering the wavelength range 1000-10000 Å. We study the impact of microturbulence and metallicity on the strength of Fe II emission and the shape of the iron emission and compare our model predictions with observational data. The primary outcomes of our study are outlined below.
@@
The unilluminated sides of the clouds do not produce enough flux to represent the observed Fe II emission
@@
A high microturbulence velocity is required to match the observed EW of Fe II and the shape of the UV Fe II emission.
3.
The updated atomic database available in the latest version of CLOUDY C23.00 could not resolve the long-standing problem of a mismatch between the expected and observed values of Fe II UV to optical ratio under the single zone approximation.
4.
@@
Appendix A Plots for the spectral fitting of RM 102 using power-law
Appendix B EW contours for microturbulence velocities 20 and 50 km s-1